AB INITIO AND EXPERIMENTAL STUDIES OF THE E INTERNAL ROTOR STATE OF He-CH_{3}F

\underline{KELLY J. HIGGINS}, ZHENHONG YU, and WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, Massachusetts 02138.

The excited E internal rotor state of He-CH_{3}F is studied using microwave spectroscopy and ab initio calculations. Bound state calculations on a high-level ab initio intermolecular potential surface^{a} predict a T-shaped ground state and two low-lying excited states with the He localized either at the C end (“linear”) or the F end (“anti-linear”) of the C-F bond, similar to that found for the A internal rotor state presented previously.^{b} There are several distinct differences between the A and E states. Whereas the ground state rotational transitions of the A state could be fit with a standard asymmetric rotor Hamiltonian, the E state exhibits a much more complex energy level structure due to the internal rotation of the CH_{3}F. In addition, the E state is calculated to be bound by 11.709 cm^{−1}, 0.249 cm^{−1}greater than the A state relative to their respective dissociation products, K = 1 or K = 0 CH_{3}F, and the ground state to linear state energy gap is predicted to be 4.148 cm^{−1}, 0.156 cm^{−1}greater than in the A state. At present six rotational transitions between five different energy levels have been observed and assigned with the aid of double resonance experiments and the ab initio calculations.

a54th International Symposium on Molecular Spectroscopy (1999)

b55th International Symposium on Molecular Spectroscopy (2000)